why torque=μNr. explain briefly.


1
Expert's answer
2021-06-27T20:44:01-0400

Gives

torque is the rotational equivalent of linear force. It is also referred to as the moment, moment of force, rotational force or turning effect, depending on the field of study. The concept originated with the studies by Archimedes of the usage of levers. Just as a linear force is a push or a pull, a torque can be thought of as a twist to an object around a specific axis. Another definition of torque is the product of the magnitude of the force and the perpendicular distance of the line of action of a force from the axis of rotation. The symbol for torque is typically τ



Torque=force"\times" perpendicular distance from the x-axis of rotation

Torque=force"\times r"

"\tau=F\times r"

"\tau=marsin\theta"

"\theta=90°,sin90°=1"

"\tau=mar"

Npartival of system

"\tau=(m_1a+m_2a+m_3a+m_4a+m_5a............+m_na)r\rightarrow(1)"

"N\mu=(m_1a+m_2a+m_3a+m_4a+m_5a............+m_na)\rightarrow(2)"

equation (1) and equation (2) we can write as

"\tau=N\mu r"